OpenStudy (anonymous):

Solve the following system of equations by graphing. x – y = 7 x – y = 5 What is the solution?

OpenStudy (across):

Graph it!

jimthompson5910 (jim_thompson5910):

No need to graph. If you let z = x-y, then you go from x-y=7 x-y=5 to z = 7 z = 5

jimthompson5910 (jim_thompson5910):

z is some (single) number...but it can't be 2 numbers at the same time

jimthompson5910 (jim_thompson5910):

So that last statement of z = 7 z = 5 is impossible